40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 24 17:12:54 2011, Koji, Update, LSC, Realignment of REFL / Some 3f PRMI locking / Recycling Gain 
    Reply  Sat Sep 24 20:14:36 2011, Koji, Update, LSC, Rough estimation of the PR gain 
    Reply  Mon Sep 26 15:15:45 2011, Anamaria, Update, LSC, Realignment of REFL / Some 3f PRMI locking / Recycling Gain 
Message ID: 5539     Entry time: Sat Sep 24 17:12:54 2011     Reply to this: 5541   5545
Author: Koji 
Type: Update 
Category: LSC 
Subject: Realignment of REFL / Some 3f PRMI locking / Recycling Gain 

[Koji Suresh]

Activity on Friday evening

- The REFL path has been thoroughly aligned
As I did not like the REFL spot misaligned on the REFL CCD, we went to the AP table.
Many optics had the spots not on the middle of the optic, including the PBS whose post was not fixed on the post holder.
We aligned the optical paths, the RF PDs, and the CCD. The alignment of the PD required the use of the IR viewer.
One should not trust the DC output as a reference of the PD alignment as it is not enough sensitive to the clipping.

We aligned the optical paths again after the reasonable alignment of PRM is established with the interferometer.
"Next time when you see REFL spot is not at the center of the camera, think what is moved!"

- The REFL165 PD is disconnected from the power supply
I found that the REFL165 PD is producing 7.5V output at the DC monitor no matter how the beam is blocked.
As I could not recover this issue by swapping the power connector at the LSC rack, I disconnected the cable
at the RFL165 PD side. I need to go through the PD power supply circuit next week.

- PRMI alignment policy of the night

Kiwamu has aligned Y-arm some time ago (Thursday evening?). I decided not to touch ITMY.
So the Michelson is aligned by ITMX, PRC is aligned by PRM.

- Michelson locking

The short Michelson was locked with AS55Q and the MICH filter. We could use the gain of +/-20 for locking,
and could increase it up to ~+/-250. At the max gain, the all three integrators and the two resonant gains
could be activated. The sign depends on which fringe you want at the AS port (bright or dark).

In this condition, the output of the POXDC channel (which is actually the POY DC out -- c.f. This entry
is used to determine the internal power. It was ~70cnt.

- PRMI locking

Then the PRMI was locked. There was some confusion of the gains because of the limitters at the servo filters
(which yielded the locking with 1bit outputs no matter how much the gains were....)

After all, I decided to use REFL33I for the PRCL for the test. The PRCL gain was -0.3~-1.0 for the carrier lock, but
was highly dependent on the alignment. i.e. if accidentally hit the high power recycling gain, it oscillated easily
and the lock was lost. Probably this was the first 3f locking at the 40m in the current optical config, if
Kiwamu did not do that secretly. The SB lock was also obtained by flipping the sign of the PRCL servo.

The difficulty we had was the instability when the recycling gain became big. We were monitoring the POXDC
(i.e.DCout of the POY PD). When this exceeds 5000, many glitches appears in the LSC signals and disturbs the lock.
This was not the fringes from neither the arms nor the SRC.

The observed POY DC with the carrier resonant PRMI was 5000~8000vcnt (momentary).

ELOG V3.1.3-